Question 2. It is a hot day, and Bert is thirsty. Here is the value he places on each bottle of water:

Value of first bottle $7

Value of second bottle $5

Value of third bottle $3

Value of fourth bottle $1

A)From this information, derive Bert’s demand schedule. Graph his demand curve for bottled water.

B)If the price of a bottle of water is $4, how many bottles does Bert buy? How much consumer surplus does Bert get from his purchases? Show Bert’s consumer surplus in your graph.


1
Expert's answer
2021-05-20T18:45:39-0400

Represent the demand Schedule as follows:




Represent the graph as follows :




If the price of the bottle of water is $4, person B buys 2 bottles (as per the demand Schedule)


Hence the Consumer Surplus ;

=Price willing to pay "-" Price actually paid


=((7+5)"-" 4×2)=$4

Consumer Surplus =$4. (AEGHI is the Consumer Surplus in the above graph)


c) Consumer Surplus =(7+5+3)"-" (3×2)=$9


(The Consumer Surplus is ABCJG)